How do you factor one side of the equation completely and find the x-intercepts and y-intercepts for y= -8x2 + x ?
Answer by kiru_khandelwal(79)

For the line with equation:
y = -8x2 + x
To find the y intercept jus put x=0 in the equation above and find the value of y.
In this case y intercept will be 0
To find x intercept put y =0 in the equation
so the equation y= -8x2 + x
Put y = 0
=> 0=-8x2 + x
=> 8x2-x = 0
=> x(8x-1) = 0
=> x= 0 or x=1/8